OpenStudy (anonymous):

Could someone please help me with this? For the following system, if you isolated y in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 5x + 3y = 9 12x + 3y = 18

jimthompson5910 (jim_thompson5910):

If you isolate y in the second equation, then... 12x + 3y = 18 3y = 18-12x 3y = -12x+18 y = -12x/3 + 18/3 y = -4x + 6 So you would substitute the expression -4x+6 in for 'y' in the first equation to go from 5x+3y = 9 to 5x+3(-4x+6) = 9

jimthompson5910 (jim_thompson5910):

From there, you can solve for x and then use that solution to find y.